OpenStudy (anonymous):

What is the solution to 15=8x2-14x

OpenStudy (anonymous):

Set it to 0 and factor

OpenStudy (anonymous):

8x^2-14x=15 8x^2-14x-15=0

OpenStudy (a1234):

Subtract fifteen from each side: 8x^2 - 14x - 15 = 0 Find two numbers that are a product of -120 (8*-15) and have a sum of -14. 8x^2 - 20x + 6x - 15 = 0 4x(2x - 5) + 3(x - 5) = 0 (4x + 3)(x - 5) = 0 4x = -3 x = -3/4 x = 5 Solutions: -3/4 and 5.